Using these data from the comparative balance sheet of Ivanhoe Company, perform horizontal analysis. (If amount and percentage are a decrease show the numbers as negative, e.g. -55,000, -20% or (55,000), (20%). Round percentages to 0 decimal places, e.g. 12%.) Increase or (Decrease) Dec. 31, 2022 Dec. 31, 2021 Amount Percentage Accounts receivable $ 362,700 $ 310,000 $Enter increase or decrease amount in dollars. Enter increase or decrease in percentages rounded to 0 decimal places. % Inventory $ 677,600 $ 560,000 $Enter increase or decrease amount in dollars. Enter increase or decrease in percentages rounded to 0 decimal points. % Total assets $3,139,500 $2,730,000 $Enter increase or decrease amount in dollars.
Answer #1
Answer

2022

2021 Amount Percentage Increase / ( Decrease )
Accounts Receiveble $        362,700 $        310,000 362700-310000 = $52,700 (52700/310000)*100 = 17% Increase
Inventory $        677,600 $        560,000 677600-560000= $ 117,600 (117600/560000)*100 = 21% Increase
Total Assets $     3,139,500 $     2,730,000 3139500-2730000= $409,500 (409500/2730000)*100 = 15% Increase

Please like